\begin{thebibliography}{10}
\bibitem{ABLP} Andries, P., Bourgat, J.-F.,  Le Tallec, P., Perthame, B.: Numerical comparison between the Boltzmann and
ES-BGK models for rarefied gases. Comput. Methods Appl. Mech. Engrg. {\bf 191} (2002), no. 31, 3369-3390.
\bibitem{ALPP} Andries,Pierre, Le Tallec, P., Perlat, J.-P., Perthame, B.: The Gaussian-BGK model of Boltzmann equation
with small Prandtl number. Eur. J. Mech. B Fluids {\bf 19} (2000), no. 6, 813-830.
\bibitem{AKT} Aoki, K., Kanba, K., Takata, S.: Numerical analysis of a supersonic rarefied gas flow past
a flat plate. Phys. Fluids, 9(4), 1997.
%\bibitem{B-G-L1} Bardos, C., Golse, F., Levermore, D.: Fluid dynamic limites of kinetic equations. I. Formal derivations,
%J. Stat. Phys. {\bf 63} (1991), no. 1-2, 323-344.
%\bibitem{B-G-L} Bardos, C., Golse, F., Levermore, D.: Fluid dynamic limites of kinetic equations. II. Convergence proofs for the Boltzmann equation,
%Comm. Pure. Appl. Math  {\bf 46} (1993), 667-753.
%\bibitem{A-G} Alonso, R.,  Gamba, I.M.: Propagation of $L^1$ and $L^{\infty}$ Maxweliian weighted bounds for derivatives of solutions
%to solutions to the homogeneous elastic Boltzmann equation. J. Math. Pures. Appl. {\bf 89}, no.6, 575-595 (2007)
%\bibitem{Ar1} Arkeryd, L.: $L^{\infty}$ estimates for the space-homogeneous Boltzmann equation, {\bf 31} no.2 347-361 (1983)
%\bibitem{Ar2} Arkeryd, L.: On the Boltzmann equation, Arch. Rational. Mech. Anal. {\bf 45} 1-34 (1972)
%\bibitem{A-E-P} Arkeryd, L. Esposito, R., Pulvirenti, M.: The Boltzmann equation for weakly inhomogeneous data. Comm. Math. Phys. {\bf 111} (3), 393-407 (1987)
%\bibitem{B-C-L} Bisi, M., Cañizo, J. A., Lods, B.: Entropy dissipation estimates for the linear Boltzmann operator. J. Funct. Anal. {\bf269} (2015), no. 4, 1028–1069.
\bibitem{BV} Berthelin, F., Vasseur, A. : From kinetic equations to multidimensional isentropic gas dynamics before shocks. SIAM J. Math. Anal. {\bf36} (2005), no. 6, 1807-1835.
\bibitem{Bello} Bellouquid, A.: Global existence and large-time behavior for BGK model for a gas with non-constant cross section, Transport Theory Statist. Phys. {\bf 32 } (2003) no. 2, 157-185.
\bibitem{BGK} Bhatnagar, P. L., Gross, E. P. and Krook, M.: A model for collision processes in gases. Small amplitude process in charged
and neutral one-component systems, Physical Revies, {\bf 94} (1954), 511-525.
\bibitem{Bird} Bird,G.A. : Molecular gas dynamics and the direct simulation of gas flows. Oxford Engineering Science, 2nd ed. (1994)
\bibitem{BCL} Bisi, M., Cañizo, J. A., Lods, B.: Entropy dissipation estimates for the linear Boltzmann operator. J. Funct. Anal. 269 (2015), no. 4, 1028–1069.
\bibitem{BC} Bobylev, A. V., Cercignani, C.: On the rate of entropy production for the Boltzmann equation. J. Statist. Phys. {\bf94} (1999), no. 3-4, 603–618
\bibitem{BoC} Bosi, R., Cáceres, M. J.: The BGK model with external confining potential: existence, long-time behaviour and time-periodic Maxwellian equilibria. J. Stat. Phys. {\bf136} (2009), no. 2, 297–330
%\bibitem{B} Boucht, F. : Construction of BGK models with a family of kinetic entropies for a given system of conservation laws. J. Statist. Phys. {\bf95} (1999), no.1-2, 113-170.
%\bibitem{BP} Bouchut, F., Perthame, B.: A BGK model for small Prandtl number in the Navier-Stokes approximation. J. Stat. Phys. {\bf 71} (1993), no. 1-2, 191-207.
%\bibitem{Duan} Duan, R.: hypocoercivity of the linearized dissipative kinetic equations, Nonlinearity, {\bf 24} (2011), no. 8, 2165-2189
%\bibitem{Duan-StrainV} Duan, R, Strain, R.: Optimal large-time behavior of the Vlasov-Maxwell-Boltzmann system in the whole space. Comm. Pure. Appl. Math. {\bf 64} (2011),no2, 375-413.
%\bibitem{Duan-STrainM} Duan, R, Strain, R.: Optimal time decay of the Vlasov-Poinsson-Boltzmann system in $\mathbb{R}^3_v$. Arch. Rational. Mech. Anal. {\bf 199} (2010), no.1, 291-328.
%\bibitem{Car}Carleman, T.: Problems mathematiques dans la theorie cineqitue des gas Uppsala: Almquist \& Wiksells 1957
\bibitem{Brull1} Brull, S.: An Ellipsoidal Statistical Model for gas mixtures, Comm. Math Sci., 13, No1, 1-13,(2015)
\bibitem{Brull2} Brull, S., Schneider, J. : A new approach of the Ellipsoidal Statistical Model. Cont. Mech. Thermodyn. 20 (2008), no.2, 63-74,
\bibitem{C} Cercignani, C.: The Boltzmann Equation and Its Application. Springer-Verlag, 1988.
\bibitem{CIP} Cercignani, C., Illner, R., Pulvirenti, M.: The Mathematical Theory of Dilute Gases. Springer-Verlag, 1994.
\bibitem{CC} Chapman, C. and Cowling, T. G.: Th emathematical theory of non-uniform gases, Cambridge University Press, 1970.
\bibitem{Chan} Chan, W. M.: An energy method for the BGK model. M. Phil thesis, City University of Hong Kong, 2007.
\bibitem{DMOS} Dolbeault, J., Markowich, P., Oelz, D., Schmeiser, C. ; Non linear diffusions as limit of kinetic equations with relaxation collision kernels. Arch. Ration. Mech, Anal. {\bf186} (2007), no.1, 133-158.
\bibitem{D-L1} DiPerna, R., Lions, P.-L.: On the Cauchy problem for the Boltzmann equation: Global exitence and weak stability. Ann. Math.
{\bf130} (2), 321-366 (1989).
%\bibitem{G-P-V} Gamba, I.M., Panferov, V., Villani, C : Upper Maxwellian Bounds for the spatially homogeneous Boltzmann equation. Arch. Rational Mech. Anal. {\bf 194} (2009) 253-282.
\bibitem{FJ} Filbet, F., Jin, S.: An asymptotic preserving scheme for the ES-BGK model of the Boltzmann equation. J. Sci. Comput. {\bf 46} (2011), no.2, 204-224.
\bibitem{FR} Filbet, F., Russo, G.: Semilagrangian schemes applied to moving boundary problems for the BGK model of rarefied gas dynamics, Kinet. Relat. Models {\bf 2} (2009), no.1, 231-250.
\bibitem{GT} Galli, M.A., Torczynski, R.: Investigation of the ellipsoidal-statistical Bhatnagar-Gross-Krook kinetic model applied to gas-phase transport of heat and tangential momentum between parallel walls, Phys. Fluids, {\bf 23} (2011) 030601
\bibitem{GL} Glassey, R.: The Cauchy Problmes in Kinetic Theory. SIAM 1996.
%\bibitem{Guo-whole} Guo, Y.: The Boltzmann equation in the whole space. Indiana Univ. Math. J. {\bf 53} (2004). no.4, 1081-1094
%\bibitem{Guo-VMB} Guo, Y.: The Vlasov-Maxwell-Boltzmann system near Maxwellians. Invent. Math. {\bf 153} (2003) no.3, 593-630
%\bibitem{Guo-VPB} Guo, Y.: The Vlasov-Poisson-Boltzmann system near Maxwellians. Comm. Pure. Appl. Math., {\bf 55} (2002) no.9, 1104-1135.
%\bibitem{Guo-CD} Guo, Y.: Continuity and decay of the Boltzmann equation in bounded domains. Arch. Ration. Mech. Anal. {\bf 197} (2010) no.3, 713-809 (2010)
%\bibitem{Guo-Strain} Guo, Y. Strain, R.: Almost exponential decay near Maxwellian, Comm. P.D.E. {\bf 31} (2006)no.2,  543-567.
%\bibitem{Guo} Guo, Y.: The Vlasov-Poisson-Boltzmann system near vacuum. Comm. Math. Phys. {\bf218} (2001), no.2, 293-313.
\bibitem{Holway} Holway, L.H.: Kinetic theory of schock structure using and ellipsoidal distribution function. Rarefied Gas Dynamics, Vol. I
(Proc. Fourth Internat. Sympos., Univ. Toronto, 1964), Academic Press, New York, (1966), pp. 193-215.
\bibitem{Issautier} Issautier, D.: Convergence of a weighted particle method for solving the Boltzmann (B.G,K.) equaiton, Siam Journal on Numerical Analysis, {\bf 33}, no 6 (1996), 2099-2199.
%\bibitem{Kawashima} Kawashima, S. : The Boltzmann equation and thirteen moments, Japan J. Appl. Math. {bf 7} (1990), 301-320.
%\bibitem{Maslova} Maslova, N.B., Tchubenko, R.P.: Dokl. Akad. Nauk. SSSR {\bf 202} (1973), 800-803. Vesin. Leningr. Univ. {\bf 1} (1973), 100-105.
%{\bf 7}, 109-113 (1976). Vestn. Leningr. Univ. {\bf 13}, 90-97 (1976)
%\bibitem{Mouhot} Mouhot, C.: Rate of convergence to equilibrium for the spatially homogeneous Boltzmann equation with hard potentials. Comm. Math. Phys. {\bf 261} no.3 629-672 (2006)
%\bibitem{L-Y-Y} Liu, T.-P, Yang, T. Yu, S.-H.: Energy method for Boltzmann equation. Phys. D {\bf 188} (2004), no. 3-4, 178-192.
\bibitem{LPS} Loyalka, S. K., Petrellis N., Storvick, T. S.: Some exact numerical results for the BGK model: Couette, Poiseuille and thermal creep
flow between parallel plates, Z. Angew. Math. Phys. 30 (3) (1979) 514–521.
\bibitem{MMM} Mellet, A., Mischler, S.,Mouhot,C. : Fractional diffusion limit for collisional kinetic equations. Arch. Ration. Mech. Anal. {\bf199} (2011),no.2 ,493-525.
\bibitem{Mischler} Mischler, S.: Uniqueness for the BGK-equation in $R^n$ and the rate of convergence for a semi-discrete scheme. Differential integral Equations {\bf 9} (1996), no.5, 1119-1138.
\bibitem{M}  Mieussens, Luc: Discrete velocity model and implicit scheme for the BGK equation of rarefied gas dynamics. Math. Models Methods Appl. Sci. {\bf10} (2000), no. 8, 1121–1149.
\bibitem{MS} Mieussens, L., Struchtrup, H. : Numerical comparison of Bhatnagar-Gross-Krook models with proper Prandtl number, Phys. Fluids {\bf 16} (2004), no.8. 2797-28
13 {\bf 10} (2000), no. 8, 1121-1149.
\bibitem{PY} Park, S., Yun, S.-B.: Cauchy problem for the ellipsoidal-BGK model of the Boltzmann equation, submitted.
\bibitem{Perthame} Perthame, B. : Global existence to the BGK model of Boltzmann equation. J. Differential Equations. {\bf 82} (1989), no.1, 191-205.
\bibitem{PPuppo} Pieraccini, S., Puppo, G.: Implicit-eplicit schemes for BGK kinetic equations. J. Sci. Comput. {\bf 32} (2007), no.1, 1-28.
\bibitem{PP} Perthame, B., Pulvirenti, M. : Weighted $L^{\infty}$ bounds and uniqueness for the Boltzmann BGK model. Arch. Rational Mech. Anal. {\bf 125} (1993), no. 3, 289-295.
\bibitem{RSY} Russo, G., Santagati, P. and Yun, S.-B. : Convergence of a semi-Lagrangian scheme for the BGK model of the Boltzmann equation. arXiv: 1007.2843v1 [math.AP].
\bibitem{SR1} Saint-Raymond, L.: From the BGK model to the Navier-Stokes equations. Ann. Sci. Ecole Norm. Sup {\bf 36} (2003), no.2, 271-317.
\bibitem{SR2} Saint-Raymond, L.: Discrete time Navier-Stokes limit for the BGK Bltzmann equation. Comm. Partial Differential Equations {\bf 27} (2002), no. 1-2, 149-184.
\bibitem{Sone} Sone, Y.: Kinetic Theory and Fluid Mechanics. Boston: Birkh\"{a}user, 2002.
\bibitem{Sone2} Sone, Y.: Molecular Gas Dynamics: Theory, Techniques, and Applications. Boston: Brikh\"{a}user, 2006.
%\bibitem{Stru} Struchtrup, H.: The BGK-model with velocity-dependent collision frequency. Contin. Mech. Thermodyn. {\bf 9} (1997), no.1 , 23-31.
\bibitem{Stru-book} Struchtrup, H.: Mesoscopic transport equaitons for rarefied gas flows: Approximation methods in kinetic theory. Springer. 2005.
\bibitem{TV} Toscani, G., Villani, C.: Sharp entropy dissipation bounds and explicit rate of trend to equilibrium for the spatially homogeneous Boltzmann equation. Comm. Math. Phys. {\bf203} (1999), no. 3, 667–706.
%\bibitem{Ukai} Ukai, S.: On the existence of global solutions of a mixed problem for the nonlinear Boltzmann equation. Proc. Japan Acad., Ser. A {\bf 53}, 179-184 (1974)
\bibitem{Ukai-BGK} Ukai, S.: Stationary solutions of the BGK model equation on a finite interval with large boundary data. Transport theory Statist. Phys. {\bf 21} (1992) no.4-6.
\bibitem{UT} Ukai, S. Yang, T.: Mathematical Theory of Boltzmann equation, Lecture Notes Series. no. 8, Liu Bie Ju Center for Math. Sci, City University of Hong Kong, 2006.
\bibitem{V} Villani, C.: A Review of mathematical topics in collisional kinetic theory. Handbook of mathematical fluid dynamics. Vol. I. North-Holland. Amsterdam, 2002, 71-305
\bibitem{V1}  Villani, C.: Cercignani's conjecture is sometimes true and always almost true. Comm. Math. Phys. 234 (2003), no. 3, 455–490.
\bibitem{Wal} Walender, P.: On the temperature jump in a rarefied gas, Ark, Fys. {\bf 7}  (1954), 507-553.
\bibitem{WZ} Wei, J., Zhang, X.: The Cauchy problem for the BGK equation with an external force.
J. Math. Anal. Appl. {\bf391} (2012), no. 1, 10–25.
\bibitem{W} Wennberg, B.: Entropy dissipation and moment production for the Boltzmann equation. J. Statist. Phys. {\bf86} (1997), no. 5-6, 1053–1066.
\bibitem{Yun} Yun, S.-B. : Cauchy problem for the Boltzmann-BGK model near a global Maxwellian. J. Math. Phy. {\bf 51} (2010), no. 12, 123514, 24pp.
\bibitem{Yun2} Yun, S.-B.: Classical solutions for the ellipsoidal BGK model with fixed collision frequency. J. Differential Equations
    {\bf259} (2015), no. 11, 6009–6037.
\bibitem{Yun3} Yun, S.-B.: Ellipsoidal BGK model near a global Maxwellian. SIAM J. Math. Anal. 47 (2015), no. 3, 2324–2354.
%\bibitem{Yun4} Yun, S.-B.: Ellipsoidal BGK model of the Boltzmann equation in the whole space
%\bibitem{Yun3} Yun, S.-B.: Ellipsoidal BGK model of the Boltzmann equation in the whole space. submitted.
\bibitem{Zhang} Zhang, X.: On the Cauchy problem of the Vlasov-Posson-BGK system: global existence of weak solutions. J. Stat. Phys. {\bf141} (2010),no.3, 566-588.
\bibitem{ZH} Zhang, X., Hu, S.: $L^p$ solutions to the Cauchy problem of the BGK equation. J. Math. Phys. {\bf 48} (2007) no.11, 113304, 17pp.
%\bibitem{Z-Stru} Zheng, Y., Struchtrup, H. : Ellipsoidal statistical Bhatnagar-Gross-Krook model with velocity dependent collision frequency. Phys. Fluids {\bf 17} (2005), 127103, 17pp.
\end{thebibliography}

\documentclass[10pt]{amsart}
\usepackage{amsmath,amsfonts, amssymb}
\usepackage[margin=1.5in]{geometry}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{corollary}{Corollary}[section]
\theoremstyle{definition}

\theoremstyle{remark}

\numberwithin{equation}{section}

\begin{document}

\title[Entropy production for ES-BGK model]{Entropy production for ellipsoidal BGK model of the Boltzmann equation}


\author{Seok-Bae Yun}
\address{Department of mathematics, Sungkyunkwan University, Suwon 440-746, Republic of Korea }
\email{sbyun01@skku.edu}



\keywords{kinetic theory of gases, BGK model, Ellipsoidal BGK model, Boltzmann equation, Entropy production}

\begin{abstract}
The ellipsoidal BGK model (ES-BGK) is a generalized version of the BGK model of the Boltzmann equation designed to yield the correct Prandtl number in the Navier-Stokes
limit. In this paper, we make two observations on the entropy production functional of the ES-BGK model. First, we show that the Cercignani type estimate holds for the ES-BGK model in the whole range of parameter $-1/2<\nu<1$.
Secondly, we observe that the ellipsoidal relaxation operator satisfies an unexpected sign-definite property.
Some implications of these observations are also discussed.
\end{abstract}

\maketitle
\section{Introduction}
The Boltzmann equation is the fundamental model bridging the particle level description and the hydrodynamic description of gases. The application of the Boltzmann equation, however, has been restrictive due mainly to  the excessive cost involved in the numerical computation of the collision integral. In this regard, Bhatnagar, Gross, Krook \cite{BGK} and,
independently, Walender \cite{Wal} suggested a relaxation type model called the BGK model. Since then, this model has been
replacing the Boltzmann equation in various numerical computations, yielding qualitatively satisfactory results
at much lower computational cost compared to that of the Boltzmann equation. There are, however, also several shortcomings reported, with the most notable one being the non-physical Prandtl number - the ratio between the viscosity and the thermal conductivity - it provides. In search of a model equation with
the correct Prandtl number, Holway \cite{Holway} derived a new equation by generalizing the local Maxwellian in the original BGK relaxation operator into
a non-isotropic Gaussian so that the stress tensor can be treated more sophisticatedly:
\begin{eqnarray}\label{ESBGK}
\begin{split}
&\partial_t f+v\cdot \nabla f=A_{\nu}(\mathcal{M}_{\nu}(f)-f),\cr
&\qquad f(x,v,0)=f_0(x,v),
\end{split}
\end{eqnarray}
which is called the ellipsoidal BGK model (ES-BGK model). $f(x,v,t)$ is the number density on the phase space of position and velocity $(x,v)\in \Omega_x\times \mathbb{R}^3$ at time $t\in \mathbb{R}^+$. Throughout this paper, $\Omega_x$
denotes $\mathbb{R}^3$ or $\mathbb{T}^3$.
The non-isotropic Gaussian $\mathcal{M}_{\nu}$ takes the following form:
\begin{eqnarray*}
\mathcal{M}_{\nu}(f)=\frac{\rho}{\sqrt{\det(2\pi\mathcal{T}_{\nu}})}\exp\left(-\frac{1}{2}(v-U)^{\top}\mathcal{T}^{-1}_{\nu}(v-U)\right),
\end{eqnarray*}
where $\rho$, $U$, $T$ and $\Theta$ are macroscopic density, bulk velocity, temperature and stress tensor $\Theta$ respectively:
\begin{eqnarray*}
\rho(x,t)&=&\int_{\mathbb{R}^3_v}f(x,v,t)dv,\cr
\rho(x,t)U(x,t)&=&\int_{\mathbb{R}^3_v}f(x,v,t)vdv,\cr
3\rho(x,t) T(x,t)&=&\int_{\mathbb{R}^3_v}f(x,v,t)|v-U(x,t)|^2dv,\cr
\rho(x,t)\Theta(x,t)&=&\int_{\mathbb{R}^3_v}f(x,v,t)(v-U)\otimes(v-U)dv,
\end{eqnarray*}
and $\mathcal{T}_{\nu}$ denotes the temperature tensor defined by
\begin{eqnarray*}
\mathcal{T}_{\nu}&=&\left(
\begin{array}{ccc}
(1-\nu) T+\nu\Theta_{11}&\nu\Theta_{12}&\nu\Theta_{13}\cr
\nu\Theta_{21}&(1-\nu)T+\nu\Theta_{22}&\nu\Theta_{23}\cr
\nu\Theta_{31}&\nu\Theta_{32}&(1-\nu) T+\nu\Theta_{33}
\end{array}
\right)\cr
&=&(1-\nu) T Id+\nu\Theta.
\end{eqnarray*}
The relaxation operator satisfies the following cancellation property:
\begin{eqnarray*}
\int_{\mathbb{R}^3_v} \big(\mathcal{M}_{\nu}(f)-f\big)\left(\begin{array}{c}1\cr v\cr|v|^2\end{array}\right)dv=0,
\end{eqnarray*}
leading to the conservations of mass, momentum and energy. H-theorem was not verified when this model was first suggested, and
proved only recently by Andries et al \cite{ALPP} (See also \cite{Brull1,Brull2}):
\[
\frac{d}{dt}\int_{\mathbb{R}^3_v}f\ln fdv\leq 0.
\]
The collision frequency $A_{\nu}$ takes the following form:
\begin{eqnarray}\label{A_nu}
A_{\nu}=\sigma(\rho,T)/(1-\nu),\quad -1/2<\nu< 1,
\end{eqnarray}
 where  $\sigma(x,y)=x^{\alpha}y^{\beta}$ for some $\alpha$, $\beta\geq0$. An application of the Chapmann-Enskog expansion shows that the Prandtl number is given by $Pr=1/(1-\nu)$ (See \cite{ALPP,CC,Holway,Stru-book}). Therefore, we can obtain the correct Prandtl number by adjusting the free parameter $\nu$. Two important cases are $\nu=(Pr-1)/Pr\approx-1/2$ and $\nu=0$.
The former gives the correct Prandtl number, while the latter corresponds to the original BGK model.\newline

The result of this paper is two-fold. First, we establish the following entropy-entropy production estimate:
\begin{eqnarray*}
D_{\nu}(f)\geq C_{\nu}H(f|\mathcal{M}_0),
\end{eqnarray*}
where $\mathcal{M}_0$ denotes the local Maxwellian constructed from $f$, and $C_{\nu}=\min\{1+2\nu,1-\nu\}A_{\nu}$.
Note that the corresponding estimate for the Boltzmann equation takes the following form \cite{TV,V1}:
\[
D_{BE}(f)\geq C_{\varepsilon}H(f|\mathcal{M}_0)^{1+\varepsilon},
\]
where $\varepsilon$ in the exponent, although it can be taken to be arbitrarily small with $C_{\epsilon}$ adjusted accordingly, can never be removed \cite{BC,W}.
Therefore, our result indicates that the ellipsoidal modification of the classical BGK model is still not sophisticate enough to capture the subtle entropy production mechanism of the Boltzmann equation: In terms of entropy production,
the BGK model behaves like the linear Boltzmann equation \cite{BCL}.
The proof is by straightforward computations making combined use of the convexity of $x\ln x$ and the concavity of $\ln x$. The following estimate comparing the entropy difference between various Maxwellians and Gaussians are crucially used:
\begin{eqnarray*}
H\big(\mathcal{M}_0\big)-H\big(\mathcal{M}_{\nu}\big)\geq \max\{\nu,-2\nu\}\big\{H\big(\mathcal{M}_0\big)-H\big(\mathcal{M}_1\big)\big\}.
\end{eqnarray*}

Secondly, we observe that the relaxation operator of the ellipsoidal BGK model has an interesting sign-definite property. We recall that, in the case of the Boltzmann equation or the original BGK model, the entropy dissipation is a
direct consequence of the following elementary inequality:
\begin{eqnarray}\label{elementary}
\mathcal{E}(a,b)\equiv(a-b)(\ln a-\ln b)\geq0.
\end{eqnarray}
The entropy production functional for the Boltzmann equation $D_{BE}(f)$ and the original BGK model $D_{0}(f)$ can be written respectively in the following manner:
\begin{align*}
D_{BE}(f)&=\int_{\mathbb{R}^3_v\times \mathbb{R}^3_{v_*}}B(v,v_*,\omega)\mathcal{E}\big(f^{\prime}f^{\prime}_*, ff_*\big)d\omega dvdv_*,\cr
D_0(f)&=\int_{\mathbb{R}^3_v}A_0\mathcal{E}\big(\mathcal{M}_0(f), f\big)dv,
\end{align*}
where $B$ denotes the Boltzmann collision kernel \cite{C,CIP,V}.
However, we immediately see that this is not the case for the ES-BGK model:
An explicit computation gives rise to an additional term $R_{\nu}(f)$ other than the one characterized by (\ref{elementary}):
\begin{align}\label{add}
D_{\nu}(f)&\equiv-\int_{\mathbb{R}^3_v}A_{\nu}\{\mathcal{M}_{\nu}(f)- f\}\ln fdv\cr
&=\int_{\mathbb{R}^3_v}A_{\nu}\mathcal{E}\big(\mathcal{M}_{\nu}(f), f\big)dv+R_{\nu}(f)
\end{align}
where the remainder term is given by
\begin{eqnarray}\label{R}
R_{\nu}( f)=\Big\{\int_{\mathbb{R}^3_v}A_{\nu}\big\{\mathcal{M}_{\nu}(f)-f\big\}(v-U)\otimes(v-U)dv\Big\}:\mathcal{T}^{-1}_{\nu}.
\end{eqnarray}
This shows that the inequality (\ref{elementary}) alone is, unlike the Boltmzann equation or the original BGK model, not sufficient to
determine the non-negativity of the entropy production functional $D_{\nu}(f)$.
The remainder term arises due to the non-conservative terms inside the ellipsoidal Gaussian, namely the stress tensor.
A rather unexpected property of this remainder $R_{\nu}$ is that it is sign definite with respect to $\nu$
in the following sense:
\[
\left\{\int_{\mathbb{R}^3_v}(\mathcal{M}_{\nu}(f)-f)(v-U)\otimes(v-U)dv\right\}:\mathcal{T}^{-1}_{\nu}\geq (\leq)~ 0\quad \mbox{ if }~\nu\geq (\leq)~0.
\]
This remainder term vanishes when $\nu=0$, which corresponds to the original BGK  model.\newline

A brief referential check is in order.
The first existence result for the BGK model goes back to  \cite{Perthame}, where the Cauchy problem were considered in the framework of
weak solutions. The uniqueness problem in a weighted $L^{\infty}$ setting  was studied in \cite{PP}.
These theories were then generalized to various directions such as the whole space case \cite{Mischler}, $L^p$ setting \cite{ZH}, and the
BGK models with external forces \cite{BoC,WZ} or mean field effect \cite{Zhang}.
Ukai studied the stationary problem in a bounded interval in \cite{Ukai-BGK}. For the results on the existence of classical solutions
and their asymptotic stability near global Maxwellians, see \cite{Bello,Chan,Yun}.
Works on various macrosopic limits such as the diffusion limit and the hydrodynamic limit can be found in
\cite{BV,DMOS,MMM,SR1,SR2}.
Literatures on the numerical computations of BGK type models are abundant, which is natural considering that it was the reason why the BGK model was first suggested.
We do not attempt to present complete list of them. Interested readers may refer to \cite{ABLP,ALPP,AKT,FJ,FR,GT,LPS,M,MS,PPuppo} and references therein.
For rigorous convergence analysis of numerical schemes, see \cite{Issautier,RSY}.

The recent revival of interest on the ES-BGK model can large be attributed to the establishment of the $H$-theorem accomplished in \cite{ALPP}.
Alternate proof was presented later by Brull et al. \cite{Brull1,Brull2}.  The mathematical study of the ES-BGK model such as the existence theory is in its initial stage. The existence of classical solutions was studied when the solution lies close to a global Maxwellian \cite{Yun3}, or when the collision frequency does not depend on macroscopic fields \cite{Yun2}.
For works on the existence of weak solutions, see \cite{PY}. Nice survey on mathematical and physical theory of kinetic equations
can be found in \cite{Bird,C,CIP,CC,GL,Sone,Sone2,Stru-book,UT,V}.
\newline

We define some notations used throughout this paper.
\begin{itemize}
\item If not stated otherwise, constants will be defined generically
\item $C_{a,b,c,..}$ denotes a generic constant which depends, not necessarily exclusively, on $a,b,c...$.
\item For $\kappa\in\mathbb{R}^3_v$, $\kappa^{\top}$ denotes its transpose.
\item Id denotes the $3\times 3$ identity matrix.
\item The Frobenius product $A:B$ denotes
\[
A:B=\sum_{1\leq i,j\leq 3}A_{ij}B_{ij}.
\]
\item We use $T$ to denote the local temperature, while $T^f$ denotes the final time.\newline
\end{itemize}

This paper is organized as follows. In section 2, we prove that the entropy production functional satisfies the Cercignani type inequality.
Exponentially fast convergence to equilibrium in the homogeneous case is derived.
In section 3, the positive definite property of the ellipsoidal relaxation operator is considered. It is shown that this property
can be used to derive the compactness in $L^1$ of the ellipsoidal Gaussian.
\newline\newline
%
%
%
\section{Entropy-Entropy production estimate for ES-BGK model}
We first define the $H$-functional $H(f)$, the relative entropy $H(f|g)$ and the entropy production functional $D_{\nu}(f)$:
\begin{eqnarray*}
H(f)&=&\int_{\mathbb{R}^3_v}f\ln fdv,\cr
H(f|g)&=&\int_{\mathbb{R}^3_v}f\ln \big(f/g\big)dv,\cr
D_{\nu}(f)&=&-\int_{\mathbb{R}^3_v}A_{\nu}\big(\mathcal{M}_{\nu}(f)-f\big)\ln fdv.
\end{eqnarray*}
Recall from (\ref{add}) that, unlike the original BGK model,
\[
D_{\nu}(f)\neq \int_{\mathbb{R}^3_v}A_{\nu}\big(\mathcal{M}_{\nu}(f)-f\big)\big(\ln \mathcal{M}_{\nu}(f)-\ln f\big)dv,
\]
except for the case $\nu=0$.
We also observe that the cases $\nu=0$ and $\nu=1$ correspond respectively to the local Maxwellian and the multivariate Gaussian with the stress tensor $\Theta$ as its covariance:
\begin{eqnarray*}
\mathcal{M}_{0}(f)&=&\frac{\rho}{\sqrt{(2\pi T)^3}}\exp\left(-\frac{|v-U|^2}{2T}\right),\cr
\mathcal{M}_{1}(f)&=&\frac{\rho}{\sqrt{\det(2\pi \Theta)}}\exp\left(-\frac{1}{2}(v-U)^{\top}\Theta^{-1}(v-U)\right).
\end{eqnarray*}
When there's no risk of confusion, we suppress the dependence on $f$ and write $\mathcal{M}_{\nu}$ instead of $\mathcal{M}_{\nu}(f)$ for simplicity.

\begin{theorem}\label{main 1} The entropy production functional $D_{\nu}$ of the ES-BGK model satisfies
\begin{eqnarray*}
D_{\nu}(f)\geq \min\{1+2\nu,1-\nu\}A_{\nu}
H(f|\mathcal{M}_{0})
\end{eqnarray*}
for $-1/2<\nu<1$.
\end{theorem}
We start with following lemma, which is crucially used in the proof of the theorem.

%
%
%
%
\begin{lemma}\label{main lemma}For $-1/2<\nu<1$
\begin{eqnarray*}
H\big(\mathcal{M}_0\big)-H\big(\mathcal{M}_{\nu}\big)\geq \max\{\nu,-2\nu\}\big\{H\big(\mathcal{M}_0\big)-H\big(\mathcal{M}_1\big)\big\}.
\end{eqnarray*}
\end{lemma}
\begin{proof}
An explicit computation gives
\begin{eqnarray*}
H(\mathcal{M}_0)&=&\int_{\mathbb{R}^3_v}\mathcal{M}_0\ln \mathcal{M}_0dv\cr
&=&\int_{\mathbb{R}^3_v}\mathcal{M}_0\Big\{\ln\frac{\rho}{(2\pi T)^{3/2}}-\frac{|v-U|^2}{2T}\Big\}dv\cr
&=&\rho\ln\frac{\rho}{(2\pi T)^{3/2}}-\frac{3}{2}\rho,
\end{eqnarray*}
and
\begin{eqnarray*}
H(\mathcal{M}_{\nu})&=&\int_{\mathbb{R}^3_v}\mathcal{M}_{\nu}\ln \mathcal{M}_{\nu} dv\cr
&=&\int_{\mathbb{R}^3_v}\mathcal{M}_{\nu}\Big\{\ln\frac{\rho}{\{\det(2\pi \mathcal{T}_{\nu})\}^{1/2}}-\frac{1}{2}(v-U)^{\top}\mathcal{T}_{\nu}^{-1}(v-U)\Big\}dv\cr
&=&\rho\ln\frac{\rho}{\{\det(2\pi \mathcal{T}_{\nu})\}^{1/2}}-\frac{1}{2}\rho\mathcal{T}_{\nu}:\mathcal{T}_{\nu}^{-1}.
\end{eqnarray*}
We then recall the identity
\begin{eqnarray*}
A:B=tr(A^{\top}B)
\end{eqnarray*}
and use the the symmetry of $\mathcal{T}_{\nu}$ to compute
\begin{eqnarray*}
\mathcal{T}_{\nu}:\mathcal{T}_{\nu}^{-1}=tr\big(\mathcal{T}^{\top}_{\nu}\mathcal{T}_{\nu}^{-1}\big)
=tr\big(\mathcal{T}_{\nu}\mathcal{T}_{\nu}^{-1}\big)=tr\big(Id\big)=3,
\end{eqnarray*}
so that
\begin{eqnarray*}
H(\mathcal{M}_{\nu}(f))=\rho\ln\frac{\rho}{\{\det(2\pi \mathcal{T}_{\nu})\}^{1/2}}-\frac{3}{2}\rho.
\end{eqnarray*}
Therefore,
\begin{eqnarray}\label{rewritten}
H(\mathcal{M}_0)-H(\mathcal{M}_{\nu}(f))=\frac{1}{2}\rho\ln\frac{\det\mathcal{T}_{\nu}}{T^3}.
\end{eqnarray}
We note that, due to the symmetry of $\Theta$, there exists an orthogonal matrix $P$ such that $P^{\top}\Theta P$ is a diagonal matrix.
We denote the diagonal elements as $\theta_i~(i=1,2,3)$, which are non-negative owing to
\[
\kappa^{\top}\Theta\kappa=\frac{1}{\rho}\int_{\mathbb{R}^3_v}f\{(v-U)\cdot\kappa\}^2dv\geq 0,\quad \kappa\in \mathbb{R}^3_v,
\]
to write
\begin{eqnarray*}
P^{\top}\Theta P=\left(\begin{array}{ccc}\theta_1&0&0\cr0&\theta_2&0\cr0&0&\theta_3\end{array}\right).
\end{eqnarray*}
This gives
\begin{eqnarray*}
P^{\top}\mathcal{T}_{\nu}P&=&P^{\top}\{(1-\nu)TId+\nu\Theta\}P\cr
&=&(1-\nu)TId+\nu P^{\top}\Theta P\cr
&=&\left(\begin{array}{ccc}(1-\nu)T+\nu\theta_1&0&0\\0&(1-\nu)T+\nu\theta_2&0\\0&0&(1-\nu)T+\nu\theta_3\end{array}\right).
\end{eqnarray*}
Since the determinant is invariant under similarity transform,
\begin{eqnarray*}
\det\mathcal{T}_{\nu}=\prod_{1\leq i\leq 3}\big\{(1-\nu)T+\nu\theta_i\big\},
\end{eqnarray*}
and (\ref{rewritten}) can be rewritten as
\begin{eqnarray}\label{rewritten2}
 H(\mathcal{M}_0)-H(\mathcal{M}_{\nu})=\frac{1}{2}\rho\sum_{i=1}^3\ln\frac{\big\{(1-\nu)T+\nu\theta_i\big\}}{T}.
\end{eqnarray}
We divide the remaining argument into the following two cases.\newline\newline
(1) $0\leq\nu<1$: From the concavity of $\ln$, we have
\begin{eqnarray*}
\ln \{(1-\nu)T+\nu\theta_i\}\geq (1-\nu)\ln T+\nu \ln \theta_i.
\end{eqnarray*}
Therefore,
\begin{eqnarray*}
 H(\mathcal{M}_0)-H(\mathcal{M}_{\nu})
&=&\frac{1}{2}\rho\left\{\sum_{i=1}^3\ln\big\{(1-\nu)T+\nu\theta_i\big\}-3\ln T\right\}\cr
&\geq&\frac{1}{2}\rho\left\{\sum_{i=1}^3\big\{(1-\nu)\ln T+\nu\ln\theta_i\big\}-3\ln T\right\}\cr
&=&\frac{1}{2}\rho\left\{\nu\sum_{i=1}^3\ln\theta_i-3\nu\ln T\right\}\cr
&=&\frac{\nu}{2}\rho\ln\frac{\theta_1\theta_2\theta_3}{T^3}\cr
&=&\nu\rho\ln\sqrt{\frac{\det \Theta}{T^{3}}}.
\end{eqnarray*}
But an explicit computation gives
\begin{eqnarray*}
\rho\ln\sqrt{\frac{\det \Theta}{T^{3}}}=H(\mathcal{M}_0)-H(\mathcal{M}_1),
\end{eqnarray*}
which yields the desired result.\newline\newline
(2) $-1/2<\nu<0$: 
In this case, $(1-\nu)T+\nu\theta_i$ is not a convex combination of $T$ and $\theta$ anymore. Instead, we observe from the similarity-invariance of the trace operator that
\begin{eqnarray*}
3T=tr(\Theta)=tr(P^{\top}\Theta P)=\theta_1+\theta_2+\theta_3
\end{eqnarray*}
to rewrite $(1-\nu)T+\nu\theta_i$ as
\begin{eqnarray*}
(1-\nu)T+\nu\theta_i&=&\frac{(1-\nu)}{3}(\theta_1+\theta_2+\theta_3)+\nu\theta_i\cr
&=&\frac{1+2\nu}{3}(\theta_1+\theta_2+\theta_3)-\nu\sum_{j\neq i}\theta_j\cr
&=&(1+2\nu)T-\nu\sum_{j\neq i}\theta_j,
\end{eqnarray*}
which is a convex combination of $T$ and $\theta_{j\neq i}$ (j=1,2,3). Now we are able to use the concavity of $ln$:
\begin{eqnarray*}
\ln\left\{(1-\nu)T+\nu\theta_i\right\}\geq (1+2\nu)\ln T-\nu\sum_{j\neq i}\ln\theta_j
\end{eqnarray*}
to proceed similarly as in the previous case:
\begin{eqnarray*}
 H(\mathcal{M}_0)-H(\mathcal{M}_{\nu})
&=&\frac{1}{2}\rho\left\{\sum_{i=1}^3\ln\big\{(1-\nu)T+\nu\theta_i\big\}-3\ln T\right\}\cr
&\geq&\frac{1}{2}\rho\left\{\sum_{i=1}^3\Big\{(1+2\nu)\ln T-\nu\sum_{j\neq i}\ln\theta_j\Big\}-3\ln T\right\}\cr
&=&-\frac{1}{2}\rho\left\{2\nu\sum_{i=1}^3\ln\theta_i-6\nu\ln T\right\}\cr
&=&-2\nu\rho\ln\sqrt{\frac{\theta_1\theta_2\theta_3}{T^3}}\cr
&=&-2\nu\rho\ln\sqrt{\frac{\det\Theta}{T^{3}}}\cr
&=&-2\nu\left\{H(M_0)-H(M_1)\right\}.
\end{eqnarray*}
Combining the case (1) and (2) gives the desired result.
\end{proof}
The following lemma says that the non-multivariate Gaussian also forms an important class of distribution functions with ``small entropy".
The proof can be found, for example, in \cite{ALPP}. We present the proof for the readers' convenience.
\begin{lemma}\cite{ALPP}\label{main lemma 2} The H-functional of $\mathcal{M}_{1}(f)$ and $\mathcal{M}_{0}(f)$ satisfies
\[
H(\mathcal{M}_{0})\leq H(\mathcal{M}_{1})\leq H(f).
\]
\end{lemma}
\begin{proof}
The first inequality is well-known. For the second one, we use the convexity of $x\ln x$, and
the fact that $\mathcal{M}_{1}(f)$ and $f$ share the same mass and the same stress tensor to compute
\begin{align*}
H(f)&\geq H(\mathcal{M}_{1})+\int_{\mathbb{R}^3_v}H^{\prime}(\mathcal{M}_{1})(f-\mathcal{M}_{1})dv\cr
&\geq H(\mathcal{M}_{1})+\int_{\mathbb{R}^3_v}\left\{1+\ln\frac{\rho}{\sqrt{\det(2\pi\Theta)}}-\frac{1}{2}(v-U)^{\top}\Theta^{-1}(v-U)\right\}\big\{f-\mathcal{M}_{1}\big\}dv\cr
&= H(\mathcal{M}_{1}).
\end{align*}
\end{proof}
\subsection{The proof of Theorem 2.1}
Due to the convexity of $x\ln x$, $D_{\nu}(f)$ satisfies
\begin{eqnarray*}
D_{\nu}(f)=-\int_{\mathbb{R}^3_v}A_{\nu}H^{\prime}(f)\big(\mathcal{M}_{\nu}(f)-f\big)dv\geq A_{\nu}\big\{H(f)-H(\mathcal{M}_{\nu})\big\}.
\end{eqnarray*}
We then split $H(f)-H\big(\mathcal{M}_{\nu}(f)\big)$ as
\begin{eqnarray*}
H(f)-H\big(\mathcal{M}_{\nu}(f)\big)&=&H(f)-H(\mathcal{M}_0)+H(\mathcal{M}_0)-H(\mathcal{M}_{\nu})\cr
&=&H(f|\mathcal{M}_0)+\left\{H(\mathcal{M}_0)-H(\mathcal{M}_{\nu})\right\},
\end{eqnarray*}
and apply Lemma \ref{main lemma} and Lemma \ref{main lemma 2} to find
\begin{eqnarray*}
H(f)-H\big(\mathcal{M}_{\nu}\big)&\geq& H(f|\mathcal{M}_0)+\max\{\nu,-2\nu\}\left\{H(\mathcal{M}_0)-H(\mathcal{M}_{1})\right\}\cr
&\geq& H(f|\mathcal{M}_0)+\max\{\nu,-2\nu\}\left\{H(\mathcal{M}_0)-H(f)\right\}\cr
&=& H(f|\mathcal{M}_0)-\max\{\nu,-2\nu\}H(f|\mathcal{M}_0)\cr
&=&\min\{1-\nu,1+2\nu\}H(f|\mathcal{M}_0).
\end{eqnarray*}
This completes the proof.

A direct consequence of Theorem 2.1, of course is the H-theorem:
\begin{corollary} H-functional for the ellipsoidal BGK model (\ref{ESBGK}) is non-increasing in time.
\end{corollary}
\begin{proof}
Multiplying $\ln f$ to (\ref{ESBGK}) and Integrating in time, we get
\begin{eqnarray}\label{corol1}
H(f(t))+\int^t_0D_{\nu}(f(s))ds=H(f_0).
\end{eqnarray}
Since $H(f|\mathcal{M}_0)\geq0$, we have from Theorem 2.1 that
\begin{eqnarray*}
D_{\nu}(f)\geq \min\{1-\nu,1+2\nu\}A_{\nu}H(f|\mathcal{M}_0)\geq0
\end{eqnarray*}
in the range $-1/2<\nu<1$, which completes the proof.
\end{proof}
\subsection{Equilibrium states are Maxwellians, not ES-Gaussians}
The relaxation operator for the ES-BGK model leaves the possibility that the non-isotropic Gaussian, not the local Maxwellian, can be an equilibrium state.
This is not a good news since it implies that the ES-BGK model may not correctly capture the asymptotic behavior of the Boltzmann equation.
Theorem 2.1, however, leads us to the conclusion that the only possible equilibrium solution is, as for the Boltzmann equation or the original BGK model, is the local Maxwellian. To show this, set
\[
\mathcal{M}_{\nu}(f)-f=0.
\]
Then, from Theorem 2.1,
\[
0=-\int_{\mathbb{R}^3_v}A_{\nu}(\mathcal{M}_{\nu}(f)-f)\ln fdv\equiv D_{\nu}(f)\geq H(f|\mathcal{M}_{0})\geq0,
\]
which implies $H(f|\mathcal{M}_{0})=0$, or
\[
f=\mathcal{M}_0=\frac{\rho}{(2\pi T)^{3/2}}\exp\left(-\frac{|v-U|^2}{2T}\right).
\]
\subsection{Exponentially fast stabilization to equilibrium in the homogeneous case}
Theorem \ref{main 1}, from a standard argument, leads to the exponential convergence to the equilibrium in the spatially homogeneous case.
Note that this is not the case for the  Boltzmann equation in general, even in the spatially homogeneous setting.
\begin{theorem} For spatially homogeneous ES-BGK model, we have
\begin{eqnarray*}
\|f(t)-\mathcal{M}_0\|_{L^1_v}\leq e^{-\frac{3}{2}\min\left\{1,\frac{1+2\nu}{1-\nu}\right\}t}
\sqrt{2H(f_0|\mathcal{M}_0)}.
\end{eqnarray*}
for $-1/2<\nu<1$.
\end{theorem}
We multiply $\ln (f/\mathcal{M}_0)$ on both sides of ES-BGK model, integrate in $v$ and apply Theorem 2.1 to get
\begin{eqnarray*}
\frac{d}{dt}H(f|\mathcal{M}_0)&\leq&-\min\{1+2\nu,1-\nu\}A_{\nu}H(f|\mathcal{M}_0)\cr
&=&-3\min\left\{1,\frac{1+2\nu}{1-\nu}\right\}H(f|\mathcal{M}_0),
\end{eqnarray*}
where we used the fact that $A_{\nu}=3/(1-\nu)$ in the homogeneous case. Therefore, by Gronwall's lemma,
\begin{eqnarray*}
H(f|\mathcal{M}_0)\leq e^{-3\min\left\{1,\frac{1+2\nu}{1-\nu}\right\}t}H(f_0|\mathcal{M}_0).
\end{eqnarray*}
We then make use of the Kullback inequality:
\begin{eqnarray*}
\|f-g\|_{L^1}\leq \sqrt{2H(f|g)}\quad \mbox{ if }\quad \int f=\int g
\end{eqnarray*}
to obtain
\begin{eqnarray*}
\|f(t)-\mathcal{M}_0\|_{L^1_v}\leq e^{-\frac{3}{2}\min\left\{1,\frac{1+2\nu}{1-\nu}\right\}t}\sqrt{2H(f_0|\mathcal{M}_0)}.
\end{eqnarray*}
%
%
%
%

%
%
%
\section{Sign-definite property of the ellipsodial relaxation operator}
In this section, we show that the relaxation operator of the ES-BGK model satisfies the following unexpected property:
\begin{theorem}\label{main 2} The remainder functional $R_{\nu}(t)$ defined in (\ref{R})
 satisfies the following sign-definiteness:
\begin{eqnarray*}
\nu R_{\nu}(x,t)\geq 0
\end{eqnarray*}
in the range $-1/2<\nu<1$. Note that $R_0(x,t)\equiv0$.
\end{theorem}
%
%
%
%
\begin{proof}
We recall the definition of $\mathcal{T}_{\nu}$ and $\Theta$ to get
\begin{eqnarray*}
\int_{\mathbb{R}^3_v}\left\{\mathcal{M}_{\nu}(f)-f\right\}(v-U)\otimes(v-U)dv
=\rho\mathcal{T}_{\nu}-\rho\Theta,
\end{eqnarray*}
and evoke the following identity again:
\begin{eqnarray*}
A:B=tr(A^{\top}B),
\end{eqnarray*}
to compute
\begin{align*}
R_{\nu}(t)&=\left\{\int_{\mathbb{R}^3_v}A_{\nu}\left\{\mathcal{M}_{\nu}(f)-f\right\}(v-U)\otimes(v-U)dv\right\}:\mathcal{T}^{-1}_{\nu}\cr
&=A_{\nu}\rho\big\{\mathcal{T}_{\nu}-\Theta\big\}:\mathcal{T}^{-1}_{\nu}\cr
&=A_{\nu}\rho\left\{tr\big(\mathcal{T}^{\top}_{\nu}\mathcal{T}^{-1}_{\nu}\big)-tr\big(\Theta^{\top}\mathcal{T}^{-1}_{\nu}\big)\right\}.
\end{align*}
Since $\mathcal{T}_{\nu}$ and $\Theta$ are symmetric matrices, $R_{\nu}(t)$ can be further simplified as
\begin{eqnarray}\label{view}
R_{\nu}(t)=A_{\nu}\rho\left\{3-tr\big(\Theta\mathcal{T}^{-1}_{\nu}\big)\right\}
\end{eqnarray}
We then recall from the proof of the Theorem \ref{main 1} that
\begin{align*}
P^{\top}\mathcal{T}^{-1}_{\nu} P&=\big\{P^{\top}\mathcal{T}_{\nu} P\big\}^{-1}\cr
&=\left(\begin{array}{ccc}
1/\{(1-\nu)T+\nu\theta_1\}&0&0\\0&1/\{(1-\nu)T+\nu\theta_2\}&0\\0&0&1/\{(1-\nu)T+\nu\theta_3\}
\end{array}\right),
\end{align*}
and use the fact that (1) $\Theta$ and $\mathcal{T}_{\nu}$ are simultaneously diagonaliziable, and (2) the trace operator is similarity invariant, to obtain the following expression for $tr\big(\Theta\mathcal{T}^{-1}_{\nu}\big)$:
\begin{align*}
tr\big(\Theta\mathcal{T}^{-1}_{\nu}\big)&=tr\left[\left(\begin{array}{ccc}\theta_1&0&0\\0&\theta_2&0\\0&0&\theta_3\end{array}\right)
\left(\begin{array}{ccc}\frac{1}{(1-\nu)T+\nu\theta_1}&0&0\\0&1\frac{1}{(1-\nu)T+\nu\theta_2}&0\\0&0&\frac{1}{(1-\nu)T+\nu\theta_3}\end{array}\right)\right]\cr
&=tr\left(\begin{array}{ccc}\frac{\theta_1}{(1-\nu)T+\nu\theta_1}&0&0\\0&1\frac{\theta_2}{(1-\nu)T+\nu\theta_2}&0\\0&0&\frac{\theta_3}{(1-\nu)T+\nu\theta_3}\end{array}\right)\cr
&=\frac{\theta_1}{(1-\nu)T+\nu\theta_1}+\frac{\theta_2}{(1-\nu)T+\nu\theta_2}+\frac{\theta_3}{(1-\nu)T+\nu\theta_3}\cr
&\equiv F_{\nu}(x,t).
\end{align*}
In view of this and (\ref{view}),
the desired estimate now follows directly once the following lemma is established.
\end{proof}
\begin{lemma}\label{F<3} $F_{\nu}$ satisfies
\begin{enumerate}
\item $F_{\nu}=3$ for $\nu=0$,
\item $F_{\nu}\leq 3$ for $0\leq \nu<1$,
\item $F_{\nu}\geq 3$ for $-1/2< \nu\leq0$.
\end{enumerate}
\end{lemma}

\begin{proof}
$(1)$ $\nu=0$: This case is simple since $F$ reduces to
\[\frac{\theta_1}{T}+\frac{\theta_2}{T}+\frac{\theta_3}{T}=3.\]
$(2)$ $0<\nu<1$: We set
\begin{eqnarray*}
A=(1-\nu)T+\nu\theta_1,\quad B=(1-\nu)T+\nu\theta_2,\quad C=(1-\nu)T+\nu\theta_3
\end{eqnarray*}
to see that
\begin{align}\label{twice0}
F_{\nu}&=\frac{\frac{1}{\nu}(A-(1-\nu)T)}{A}+\frac{\frac{1}{\nu}(B-(1-\nu)T)}{B}+\frac{\frac{1}{\nu}(C-(1-\nu)T)}{C}\cr
&=\frac{1}{\nu}\left\{3-(1-\nu)T\left(\frac{1}{A}+\frac{1}{B}+\frac{1}{C}\right)\right\}
\end{align}
Applying the arithmetic inequality twice, we obtain
\begin{eqnarray}\label{twice}
\frac{1}{A}+\frac{1}{B}+\frac{1}{C}\geq\frac{3}{\sqrt[3]{ABC}}\geq \frac{3}{\frac{A+B+C}{3}}=\frac{3}{T}.
\end{eqnarray}
In the last part, we used
\begin{eqnarray*}
\sum\theta_i=tr(P^{\top}\Theta P)=tr(\Theta)=3T,
\end{eqnarray*}
and
\begin{eqnarray*}
A+B+C
=3(1-\nu)T+\nu(\theta_1+\theta_2+\theta_3)=3T.
\end{eqnarray*}
Therefore, from (\ref{twice0}) and (\ref{twice}), we get the desired result:
\begin{eqnarray*}
F_{\nu}\leq\frac{3}{\nu}-\frac{(1-\nu)T}{\nu}\left(\frac{3}{T}\right)\leq 3.
\end{eqnarray*}
$(3)$ $-1/2<\nu<0$: This case follows in the exactly same manner as in $(2)$. The only difference is the
change of direction of the inequality, due to the negative sign of $\nu$.
\end{proof}

%
%
%
%
%
\subsection{Sign-definite property at linearized level} The sign-definite property of the remainder term in entropy production functional can be understood in a more explicit way in the linearized setting.
Let $f=m+\sqrt{m}g$ for $m=1/(2\pi)^{3/2}e^{-|v|^2/2}$, then $(\ref{ESBGK})$ is rewritten as
\[
\partial_tg+v\cdot\nabla_xg=L_{\nu}g+\Gamma(g).
\]
The linearized relaxation operator $L_{\nu}$  takes the following form:
\[
L_{\nu}g=\frac{1}{1-\nu}\left\{(P_0f-f)+\nu (P_1f+P_2f)\right\}.
\]
For the definition of $\Gamma(g)$, which is not relevant for our purpose, see \cite{Yun2}.
$P_0$, $P_1$ and $P_2$ denote the projection operators on the linear spaces spanned respectively by
\[
\{1,v,|v|^2\}, \quad\{3v_i^2-|v|^2\}_{i=1,2,3},~\mbox{ and }~\{v_iv_j\}_{i<j}.
\]
It was shown in \cite{Yun2} that  $P_0\perp(P_1+P_2)$, which gives
\begin{eqnarray*}
-\langle L_{\nu}f,f\rangle_{L^2_v}=\frac{1}{1-\nu}\|(I-P_0)f\|^2_{L^2_v}+\frac{\nu}{1-\nu}\|(P_1+P_2)f\|^2_{L^2_v}.
\end{eqnarray*}
This shows that the linearized entropy dissipation $-\langle L_{\nu}f,f\rangle_{L^2_v}$ is decomposed into the usual dissipation term
and the additional remainder not present in the original BGK model or the Boltzmann equation.
The remainder term $\nu(1-\nu)^{-1}\|(P_1+P_2)f\|^2_{L^2_v}$ clearly is sign definite with respect to $\nu$ in the range $-1/2<\nu<1$.
%
%
%
%
%
\subsection{Weak compactness of $\mathcal{M}_{\nu}(f)$ in $L^1$}
Theorem \ref{main 2} can be employed to show the compactness of the ellipsoidal Gaussian by enabling one to derive the Diperna-Lions type inequality \cite{D-L1} in the range $-1/2<\nu<1$, when $A_{\nu}$ does not depend on macroscopic fields. We first consider
\begin{align*}
\mathcal{M}_{\nu}(f)-f&=\big\{\mathcal{M}_{\nu}(f)-f\big\}\big(1_{\mathcal{M}_{\nu}(f)<Rf^n}+1_{\mathcal{M}_{\nu}(f)>Rf}\big)\cr
&\leq(R-1)f1_{\mathcal{M}_{\nu}(f)<Rf}\cr
&+\frac{1}{\ln R}\big(\mathcal{M}_{\nu}(f)-f\big)\big(\ln\mathcal{M}_{\nu}(f)-\ln f\big)1_{\mathcal{M}_{\nu}(f)\geq Rf},
\end{align*}
where $R>1$.
This gives, for any fixed positive number $T^f$ and measurable set $B_{x,v}\subset\Omega_x\times\mathbb{R}^3_v$,
\begin{eqnarray*}
&&\int^{T^f}_0\int_{B_{x,v}}\mathcal{M}_{\nu}(f)dxdvdt\cr
&&\quad\leq R\int^{T^f}_0\int_{B_{x,v}}fdxdvdt+\frac{1}{\ln R}\int^{T^f}_0\int_{ B_{x,v}}\big(\mathcal{M}_{\nu}(f)-f\big)\big(\ln\mathcal{M}_{\nu}(f)-\ln f\big)dvdxdt\cr
&&\quad\leq R\int^{T^f}_0\int_{B_{x,v}}fdxdvdt+\frac{1}{\ln R}\int^{T^f}_0\int_{ \Omega_x\times \mathbb{R}^3_v}\big(\mathcal{M}_{\nu}(f)-f\big)\big(\ln\mathcal{M}_{\nu}(f)-\ln f\big)dvdxdt\cr
&&\quad= R\int^{T^f}_0\int_{B_{x,v}}fdxdvdt
+\frac{1}{\ln R}\int^{T^f}_0\int_{ \Omega_x}D_{\nu}(f)dxdt-\frac{1}{\ln R}\int^{T^f}_0\int_{\Omega_x}R_{\nu}(f)dxdt\cr
&&\quad\leq R\int^{T^f}_0\int_{B_{x,v}} fdxdvdt+\frac{1}{\ln R}\left\{\int_{\Omega_x\times\mathbb{R}^3_v}f_0|\ln f_0|dxdv+C_{f_0,T^f}\right\}\cr
&&\quad-\frac{1}{\ln R}\int^{T^f}_0\int_{\Omega_x}R_{\nu}(f)dxdt.
\end{eqnarray*}
Here we used
\[
\int_{\Omega_x\times \mathbb{R}^3_v}f(t)|\ln f(t)|dxdv+\int^t_0\int_{\Omega_x}D_{\nu}(f(s))dxds\leq
\int_{\Omega_x\times \mathbb{R}^3_v}f_0|\ln f_0|dxdv+C_{f_0,t},
\]
which follows by a standard argument from (\ref{corol1})\cite{D-L1}.
When $0<\nu<1$, the last term is non-negative, and can be ignored:
\begin{align*}
\int^{T^f}_0\int_{B_{x,v}}\mathcal{M}_{\nu}(f)dxdvdt&\leq
R\int^{T^f}_0\int_{B_{x,v}} fdxdvdt\cr&+\frac{1}{\ln R}\left\{\int_{\Omega_x\times \mathbb{R}^3_v}f_0|\ln f_0|dxdv+C_{f_0,T^f}\right\}.
\end{align*}
In the case $-1/2<\nu\leq 0$, we recall
\[
R_{\nu}=A_{\nu}\rho(3-F_{\nu}),
\]
and
\[
(1-\nu)T+\nu\theta_i=(1+2\nu)T-\nu\sum_{j\neq i}\theta_j,
\]
to compute
\begin{eqnarray*}
&&-\int^{T^f}_0\int_{\Omega_x}R_{\nu}(x,t)dxdt\cr
&&\hspace{0.9cm}=\int^{T^f}_0\int_{\Omega_x}A_{\nu}\rho\left\{-3+\sum_{1\leq i\leq 3}\frac{\theta_i}{(1-\nu)T+\nu\theta_i}\right\}dxdt\cr
&&\hspace{0.9cm}=-3A_{\nu}\int^{T^f}_0\int_{\Omega_x\times \mathbb{R}^3_v}fdxdvdt+A_{\nu}\sum_{1\leq i\leq 3}\int^{T^f}_0\int_{ \Omega_x}\frac{\theta_i\rho}{(1-\nu)T+\nu\theta_i} dxdt\cr
&&\hspace{0.9cm}=-3A_{\nu}\int^{T^f}_0\int_{\Omega_x\times \mathbb{R}^3_v}fdxdvdt+A_{\nu}\sum_{1\leq i\leq 3}\int^{T^f}_0\int_{\Omega_x}\frac{\theta_i\rho}{(1+2\nu)T-\nu\sum_{j\neq i}\theta_j} dxdt\cr
&&\hspace{0.9cm}\leq-3A_{\nu}\int^{T^f}_0\int_{\Omega_x\times \mathbb{R}^3_v}fdxdvdt+A_{\nu}\sum_{1\leq i\leq 3}\int^{T^f}_0\int_{\Omega_x}\frac{\theta_i\rho}{(1+2\nu)T} dxdt\cr
&&\hspace{0.9cm}\leq-3A_{\nu}\int^{T^f}_0\int_{\Omega_x\times \mathbb{R}^3_v}fdxdvdt+\frac{3A_{\nu}}{1+2\nu}\int^{T^f}_0\int_{\Omega_x}\rho dxdt\cr
&&\hspace{0.9cm}\leq\frac{-6\nu}{(1-\nu)(1+2\nu)}\int^{T^f}_0\int_{\Omega_x\times \mathbb{R}^3_v}fdxdvdt\cr
&&\hspace{0.9cm}=\frac{-6\nu T_f}{(1-\nu)(1+2\nu)}\int_{\Omega_x\times \mathbb{R}^3_v}f_0dxdv.
\end{eqnarray*}
In summary, we have
\begin{align*}
\int^{T^f}_0\int_{B_{x,v}}\mathcal{M}_{\nu}(f)dxdvdt&\leq R\int^{T^f}_0\int_{B_{x,v}} fdxdvdt\cr
&+\frac{1}{\ln R}\left\{\int_{\Omega_x\times\mathbb{R}^3_v}f_0|\ln f_0|dxdv+C_{f_0,T^f}\right\}\cr
&+\frac{1}{\ln R}\left\{\frac{-6\nu T_f}{(1-\nu)(1+2\nu)}\int_{\Omega_x\times \mathbb{R}^3_v}f_0dxdv\right\}1_{-1/2<\nu<1}.
\end{align*}
which, from the Dunford-Pettis theroem, gives the weak compactness of $\mathcal{M}_{\nu}(f)$ in $L^1((0,T)\times \Omega_x\times \mathbb{R}^3_v)$ once
$\{f\}$ is weak compact in $L^1((0,T)\times \Omega_x\times \mathbb{R}^3_v)$.
\begin{section}{Acknowledgement}
This research was supported by Basic Science Research Program through
the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT $\&$
Future Planning (NRF-2014R1A1A1006432)
\end{section}


%
%
\bibliographystyle{amsplain}

\end{document}